A simple machine with an efficiency of 75% lift a load of 5000N. When a Force of 500N is applied to it, Calculate the velocity of ratio of the machine
1
Expert's answer
2020-12-06T17:15:25-0500

The mechanical advantage


"M.A.=\frac{L}{E}=\frac{5000}{500}=10"


The velocity ratio


"V.R.=\frac{M.A.}{\eta}=\frac{10}{0.75}=13.3"